An object moves such that its position on the x-axis is given by the function x(t)=a+bt+ct2, where t is time. Given that x(1)=10, x(2)=11, and x(3)=2 find the coefficients a, b, and c.

(Linear Algebra Question)
Answer #1

Given that x(t)= a + b + ct2 and x (4) = 10, X62) - 11, (x(3)=2. Given x ct) = at bttct² put x c1)=10, le tot in -cij So, a+b